OpenStudy (anonymous):

Divide & Simplify b/b^-1

OpenStudy (anonymous):

\[\frac{b}{b^{-1}}=b\times b=b^2\]

OpenStudy (anonymous):

b/b^-1 = b * b =b^2

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

\[\frac{a_1}{a_2} = a_1\cdot a_2^{-1}\]So in our case: \[\frac{b}{b^{-1}} = b \cdot {b^{-1}}^{-1} = b \cdot b = b^2 \]